The following sample data are from a normal population: 14, 12, 16, 19, 17, 15, 10, 9.

(a) What is the point estimate of the population mean?

(b) What is the point estimate of the population standard deviation? (Round your answer to three decimal places.)

(c) With 95% confidence, what is the margin of error for the estimation of the population mean? (Round your answer to one decimal place.)

(d) What is the 95% confidence interval for the population mean? (Round your answer to one decimal place.) to
Answer #1

a)

sample mean, xbar = 14

b)


sample standard deviation, s = 3.464

C)


sample size, n = 8
degrees of freedom, df = n - 1 = 7

Given CI level is 95%, hence α = 1 - 0.95 = 0.05
α/2 = 0.05/2 = 0.025, tc = t(α/2, df) = 2.365


ME = tc * s/sqrt(n)
ME = 2.365 * 3.464/sqrt(8)
ME = 2.9

d)

CI = (xbar - tc * s/sqrt(n) , xbar + tc * s/sqrt(n))
CI = (14 - 2.365 * 3.464/sqrt(8) , 14 + 2.365 * 3.464/sqrt(8))
CI = (11.1 , 16.9)